please help me solve this equation : 2x/x+2=1
Found 2 solutions by satyareddy22, josgarithmetic:
Answer by satyareddy22(84)

2x/x+2=1
2x=1(x+2)
2x=x+2
2x-x=2
x=2

Answer by josgarithmetic(39617)

The equation does not work as written. Try it this way: 2x/(x+2)=1.


x=2